Time Warner shares have a market capitalization of $70 billion. The company is expected to pay a dividend of $0.25 per share and each share trades for $30. The growth rate in dividends is expected to be 8​%per year. ​ Also, Time Warner has $20 billion of debt that trades with a yield to maturity of 7​%. If the​ firm's tax rate is 35​%, compute the​ WACC?  

A. 7.49​%

B. 6.7​%

C. 7.88​%

D. 7.09​%

Please show your work on excel
Answer #1

ANSWER IS C 7.88%

here is excel solution for this step by step

using dividend discount model share price = dividend expected/ (cost of capital - growth rate)
30= 0.25/(c-8%)
c-8%= 0.25/30
c-8%= 0.00833
c= 0.088333333
cost of equity = 8.83%
cost of debt = YTM(1-tax rate)
= 7% x (1-35%)
cost of debt = 4.55%
total capital= 70+20= 90
WACC = (weight of equity x cost of equity )+ (weight of debt x cost of debt)
WACC = ((70/90) x 8.83%) + ((20/90) x 4.55%)
WACC = 0.078788889
WACC = 7.88%